# Leveraging Quantum Superposition to Infer the Dynamic Behavior of a Spatial-Temporal Neural Network Signaling Model

### v2

@@ 3.3.3 Internal dynamics of $v_{j}$
Let $f_{j}(\Omega_{j},t)$ represent the instantaneous state of a vertex $v_{j}$ as a function of time and an internal node model or activation function with parameter set $\Omega_{j}$. Informally, the internal state can be interpreted as a binary function, which we define at an observation time $T_{o}$ as:
$f_{j}(\Omega_{j},T_{o})=\begin{cases}1,&\text{if $v_{j}$ can respond to an input,}\\ 0,&\text{if it is refractory to any input.}\end{cases}$ (1)
In the simpler single activation case, we assume that the first arriving discrete signal from any of the $v_{i}$ vertices connected into $v_{j}$ can activate it. Once this occurs, $v_{j}$ becomes refractory for a duration $R_{j}$, during which $f_{j}=0$. Below, we will progressively build out more neurobiologically realistic conditions (that do not rely on just one signaling event activating $v_{j}$), and we will formalize what we mean by ’can respond to an input’ and ’is refractory to an input’.
@@ 3.3.4 Refraction ratio and signal competition
To compute, in parallel, which vertex $v_{i}\in H_{j}(v_{i})$ causally activates $v_{j}$ at discrete times, i.e., $H_{j}[v_{i}]\leadsto v_{j}$ in the notation we established above, we track the ’position’ of propagating signals along edges, and therefore the amount of latency before they reach target vertices, relative to the refractory states of vertices. This is achieved using the refraction ratio, defined as the ratio between the refractory period $R_{j}$ and the signaling latency $\tau_{ij}$ associated with a discrete signal from vertex $v_{i}$ to $v_{j}$:
@@
This ratio provides a basis for computing and determining $H_{j}[v_{i}]\leadsto v_{j}$ by analyzing $\Delta_{ij}$ for all $v_{i}\in H_{j}(v_{i})$.
@@ 3.3.5 Bounds and constraints on $\Delta_{ij}$
    As $R_{j}\to 0$, $y_{j}=1$ at all times, implying an instantaneous recovery from refractoriness, which is physically unrealizable. Hence, $R_{j}>0$.
@@
    As $R_{j}\to\infty$, $y_{j}=0$ at all times, resulting in no information flow or signaling. Thus, $0<R_{j}<\infty$.
@@ 3.3.6 Effective refractory period and temporal offsets
If $v_{j}$ becomes refractory before an arbritrary observation time $T_{o}$, it could be in this state for a residual time less than $R_{j}$ at $T_{o}$. For example, if $v_{j}$ is halfway recovered from its refractory period at the time the node is observed at $T_{o}$, then the remaining amount of time before it can be activated again will be $R_{j}/2$, and not the full period $R_{j}$. This requires accounting for ’how much time is left’ in both the signaling latencies $\tau_{ij}$ and the refractory periods $R_{j}$ at $T_{o}$ to compute activation patterns. We define the residual refractory period at $T_{o}$ as the effective refractory period, denoted $\bar{R}_{j}$, which is central to our model.
@@ 3.5.1 Probabilistic extension
As developed in the preceding sections, the framework is deterministic, meaning that a single vertex $v_{i}$ is guaranteed to activate $v_{j}$ when the condition $H_{j}[v_{i}]\leadsto v_{j}$ is satisfied. In other words, the probability of $v_{j}$ responding to a winning signal from $v_{i}$ is exactly 1. However, biological neural networks are not deterministically precise and are often influenced by random fluctuations in signaling processes. These fluctuations may arise due to thermal dynamics, sub-diffusion processes, or molecular stochasticity, such as neurotransmitter vesicles crossing the synaptic cleft or the probabilistic binding of neurotransmitters on the postsynaptic membrane.
We define a probability distribution $P_{ij}$ to represent the likelihood that a winning vertex $H_{j}[v_{i}]\leadsto v_{j}$ successfully activates $v_{j}$. Let $v_{j}(P_{ij})$ denote the output probability of $v_{j}$ for a given threshold $P_{threshold}$. The activation condition for $v_{j}$ based on $P_{ij}$ can then be defined as:
@@ 3.5.3 Fractional summation of contributing signaling events
With the model now fully developed, we can revisit the condition for the state of node $v_{j}$, given by $f_{j}$, as informally introduced in equation 1. The state of $v_{j}$ at an arbitrary observation time $T_{o}$ is given by
$F(V,T_{o})\coloneqq f_{j}(T_{o})=\begin{cases}&\text{if}\;\Sigma_{r}\geq\Sigma_{T},\;v_{j}=1\\ &\text{if}\;\Sigma_{r}<\Sigma_{T},\;v_{j}=0\\ \end{cases}$ (16)
Determining the state given by equation 15 for every node in the network by computing equation 15 in parallel at an observation or measurement time $T_{o}$ defines the network dynamical state at $T_{o}$.
Assuming an excitatory network, i.e., a network where for all nodes if a node is activated it generates an excitatory output, we are now in a position to formally restate the key question we ask in this paper: Given some (observed or unobserved) period of evolution dynamics of a network computed by equation 15, can the dynamics (state) of the network sustain inherent activity beyond an arbitrary observation time $T_{o}$, or are the dynamics guaranteed to stop, either due to ’epileptic’-like saturation or quiescence? The (global) state of the network is determined by $F(V,T_{o})=f_{j}(T_{o})\forall v\in V$ (equation 16).
@@ 4 Preparing Network States for Deutsch-Jozsa Computations
To leverage the Deutsch–Jozsa algorithm, we need to structure the problem so that it naturally exploits the algorithm’s deterministic superposition properties. This will allow us to determine the global state of the network all at once, rather than needing to classically compute and check the state of every $v_{j}$ individually.
To achieve this, we will first construct a specific set that encodes the states of $v_{j}\in V$. When presented as an input to Deutsch–Jozsa, the output can be interpreted as an answer to the question about the network’s dynamics. We will also use Grover’s algorithm to evaluate set membership within this framework.
@@ 4.1 Set Membership Conditions
Each vertex $v_{j}\in V$ is associated with a value $\Sigma_{r}\in\mathbb{R}$ at time $T_{o}$, that is typically normalized to a defined range:
@@
In this context, $F(V,T_{o})$ is a function that compares the $n$-bit input $\Sigma_{r}$, the running summation for each vertex $v_{j}\in V$, to a constant threshold $\Sigma_{T}\in\{0,1\}^{n}$ (equation 16). To determine the global state of the network, we evaluate $F(V,T_{o})$ for all $v_{j}\in V$ at time $T_{o}$, where the output is either 00 or $1$ for each node, indicating whether $v_{j}$ fires (activates) at the next time step (c.f. Section 3.5.3).
@@
Mathematically, we construct an ordered list $\mathbb{V}$ such that its elements map each vertex $v_{j}\in V$ to its corresponding value of $\Sigma_{r}$:
$\mathbb{V}:v_{j}\rightarrow\Sigma_{r}\quad\forall v_{j}\in V.$ (17)
@@
    $S_{n}\coloneqq\{s_{i}\,|\,s_{i}\in\mathbb{S}_{n}\,\wedge\,s_{i}=\Sigma_{r}\in\mathbb{V}\}.$ (19)
@@
    $S_{n}^{c}\coloneqq\{s_{i}\,|\,s_{i}\in\mathbb{S}_{n}\,\wedge\,s_{i}\neq\Sigma_{r}\in\mathbb{V}\}.$ (20)
@@
There exists a surjective mapping $\mathbb{V}\rightarrow\mathbb{S}_{n}$: Each $v_{j}\in V$ maps to one of the $n$-bit string values $s_{i}\in\mathbb{S}_{n}$. In general, we expect:
$|\mathbb{V}|=|V|=N>|\mathbb{S}_{n}|\geq|S_{n}|.$
@@
Given this construction, another canonical quantum computing algorithm, Grover’s algorithm, can determine set membership in $S_{n}$. The interpretation of $S_{n}$ in the context of how the question about neural dynamics is framed motivates the construction of the unitary function to be evaluated by the Deutsch–Jozsa algorithm later on.
4.2 Functional Interpretations of $F(V,T_{o})$ Outcomes
Recall that our goal is to determine if the network’s dynamics can sustain inherent activity beyond $T_{o}$, or whether it is guaranteed to stop due to either ’epileptic’ saturation or quiescence. This is achieved by evaluating the global state of the network using $f_{j}(T_{o})$.
@@ 4.2 Functional Interpretations of $F(V,T_{o})$ Outcomes
    $F(V,T_{o})=f_{j}(T_{o})=0\quad\forall v_{j}\in V$
In both cases, $F(V,T_{o})$ is constant (in the language of Deutsch–Jozsa), mapping $S_{n}$ to binary outputs $\{0,1\}$. In contrast, inherent dynamic activity will occur only when $F(V,T_{o})$ is not constant, which we can interpret as the network being able to sustain autonomous activity for at least one additional time step.
By running the Deutsch–Jozsa algorithm twice, we can distinguish between constant and non-constant states of $F(V,T_{o})$. A non-constant result implies that the network is neither epileptic nor quiescent. It is important to note that $F(V,T_{o})$ need not be balanced — only not constant. This agrees with the Deutsch–Jozsa requirement for unitary functions of the form $f(\{0,1\}^{n})\rightarrow\{0,1\}$.
It is critical to note in the statement above that this is a special use case of the Deutsch-Jozsa algorithm. Deutsch-Jozsa takes as input a function that evaluates a binary sequence of 0’s and 1’s, i.e., a set $\{0,1\}^{n}$, and uses quantum superposition to simultaneously determine if the input function’s output is constant (all 0’s or all 1’s) or whether the output is balanced (an equal number of 0’s and 1’s). To evaluate $F(V,T_{o})=F(S_{n})$ using Deutsch-Jozsa, we first need to extend $S_{n}$ to a construction compatible with the algorithm. In our case, we need to evaluate $S_{n}$ as inputs and not the full domain $\mathbb{S}_{n}=\{0,1\}^{n}$. Deutsch-Jozsa assumes that the oracle function being evaluated is defined over the entire domain $\mathbb{S}_{n}$. It cannot evaluate a particular subset of $\{0,1\}^{n}$. Thus, it cannot explicitly evaluate $F(V,T_{o})=F(S_{n})$. We show how to carry this out below.
@@ 4.3 Determining Membership in $S_{n}$ and $S_{n}^{c}$ Using Grover’s Algorithm
To identify which elements of $\mathbb{S}_{n}$ belong to $S_{n}$, we need to search for each $n$-bit string $s_{i}\in\mathbb{S}_{n}$ that is in the list $\mathbb{V}$, which contains the values of $\Sigma_{r}$ resulting from the dynamic evolution of the network up to the observation time $T_{o}$. The resultant set is $S_{n}$, and its complement in $\mathbb{S}_{n}$ forms $S_{n}^{c}$. Recall that we expect some elements of $\mathbb{S}_{n}$ to appear in $\mathbb{V}$ only once, some multiple times, and others not at all. Grover’s algorithm offers a potential quadratic speedup over classical approaches to construct $S_{n}$ by enabling an efficient search within $\mathbb{S}_{n}$.
@@ 4.3.1 Construction of the oracle function
The list $\mathbb{V}$ contains the running summation values $\Sigma_{r}$ for each vertex $v_{j}$ at $T_{o}$ (see equation 17). Assume each element in $\mathbb{V}$ is labeled as $\{\nu_{1},\nu_{2},\dots,\nu_{M}\}$. Each of these elements of $\mathbb{V}$ corresponds to an $n$-bit string in $\mathbb{S}_{n}$. For each $s_{i}\in\mathbb{S}_{n}$, we construct an oracle function $U_{f}$ to identify if $s_{i}\in\mathbb{V}$. Specifically, $U_{f}$ checks if there exists a $\nu_{i}\in\mathbb{V}$ such that $\nu_{i}=s_{i}$ for some $s_{i}\in\mathbb{S}_{n}$:
$\exists\nu_{i}\in\mathbb{V}\text{ such that }\nu_{i}=s_{i}.$ (21)
@@
where |$\ket{\nu_{i}}$ encodes a state corresponding to $\nu_{i}\in\mathbb{V}$, and $f(\nu_{i})$ evaluates to 1 if $\nu_{i}=s_{i}$, and 0 otherwise.
To construct $U_{f}$, a programmable oracle function will be required on a physical machine capable of implementing the logical condition ‘if the current state $\nu_{i}=s_{i}$, apply a phase flip; otherwise, do nothing’. This could be achieved through conditional quantum gates for each $s_{i}$. The technical and hardware implementations to achieve this carry their own challenges. Our focus in this paper, however, is on the theoretical and algorithmic constructions and considerations of the problem setup and solution. We leave a discussion about the possible approaches to the physical computation of the theoretical structures we derive in this section and the rest of the paper for future efforts.

### v4

@@ 3.3.3 Internal dynamics of $v_{j}$
Let $\mathcal{T}(v_{j})$ represent the instantaneous state of a vertex $v_{j}$ as a function of some internal node model or activation function at an arbitrary observation or measurement time $T_{o}$. Informally, the internal state can be interpreted as a binary function, which we define as:
$\mathcal{T}(V)=\begin{cases}1,&\text{if $v_{j}$ can respond to an input,}\\ 0,&\text{if it is refractory to any input.}\end{cases}$ (1)
Note that we wrote $\mathcal{T}(V)$ in equation 1 as shorthand to represent the operation of $\mathcal{T}(\cdot)$ on all $v_{j}\in V$, but will write $\mathcal{T}(v_{j})$ to indicate its operation on the specific vertex $v_{j}$. In the simpler single activation case, we assume that the first arriving discrete signal from any of the $v_{i}$ vertices connected into $v_{j}$ can activate it. Once this occurs, $v_{j}$ becomes refractory for a duration $R_{j}$, during which $\mathcal{T}(v_{j})=0$. Below, we will progressively build out more neurobiologically realistic conditions (that do not rely on just one signaling event activating $v_{j}$), and we will formalize what we mean by ’can respond to an input’ and ’is refractory to an input’.
@@ 3.3.4 Refraction ratio and signal competition
To compute (in parallel for all vertices) which vertex $v_{i}\in H_{j}(v_{i})$ causally activates $v_{j}$ at discrete times, i.e., $H_{j}[v_{i}]\leadsto v_{j}$ in the notation we established above, we track the ’position’ of propagating signals along edges, and therefore the amount of latency before they reach target vertices, relative to the refractory states of the vertices. This is achieved using the refraction ratio, defined as the ratio between the refractory period $R_{j}$ and the signaling latency $\tau_{ij}$ associated with a discrete signal from vertex $v_{i}$ to $v_{j}$:
@@
This ratio provides a basis for computing and determining $H_{j}[v_{i}]\leadsto v_{j}$ by analyzing $\Delta_{ij}$ for all $v_{i}\in H_{j}(v_{i})$ at any given moment.
@@ 3.3.5 Bounds and constraints on $\Delta_{ij}$
    As $R_{j}\to 0$, $f_{j}=1$ at all times, implying an instantaneous recovery from refractoriness, which is physically unrealizable. Hence, $R_{j}>0$.
@@
    As $R_{j}\to\infty$, $f_{j}=0$ at all times, resulting in no information flow or signaling. Thus, $0<R_{j}<\infty$.
@@ 3.3.6 Effective refractory period and temporal offsets
If $v_{j}$ becomes refractory before an arbritrary observation time $T_{o}$, it could be in this state for a residual time less than the full recovery period $R_{j}$ at $T_{o}$. For example, if $v_{j}$ is halfway recovered from its refractory period at the time the node is observed at $T_{o}$, then the remaining amount of time before it can be activated again will be $R_{j}/2$, and not the full period $R_{j}$. This requires accounting for ‘how much time is left’ in both the signaling latencies $\tau_{ij}$ and the refractory periods $R_{j}$ at $T_{o}$ to compute activation patterns. We define the residual refractory period at $T_{o}$ as the effective refractory period, denoted $\bar{R}_{j}$, which is central to our model.
@@ 3.5.1 Probabilistic extension
As developed in the preceding sections, the framework is deterministic, meaning that a single vertex $v_{i}$ is guaranteed to activate $v_{j}$ when the condition $H_{j}[v_{i}]\leadsto v_{j}$ is satisfied. In other words, the probability of $v_{j}$ responding to a winning signal from $v_{i}$ is exactly 1. However, biological neural networks are not deterministically precise and are often influenced by random fluctuations in molecular and cellular signaling processes. These fluctuations may arise due to thermal dynamics, sub-diffusion processes, or molecular stochasticity, such as neurotransmitter vesicles crossing the synaptic cleft or the probabilistic binding of neurotransmitters on the postsynaptic membrane.
We define a probability distribution $P_{ij}$ to represent the likelihood that a winning vertex $H_{j}[v_{i}]\leadsto v_{j}$ successfully activates $v_{j}$. Let $v_{j}(P_{ij})$ denote the probability of an effect on $v_{j}$, such as its activation (but see the below), for a given threshold $P_{threshold}$. The activation condition for $v_{j}$ based on $P_{ij}$ can then be defined as:
@@ 3.5.3 Fractional summation of contributing signaling events
With the model now fully developed, we can revisit the condition for the state of node $v_{j}$, given by $\mathcal{T}(V)$, as informally introduced in equation 1. The state of $v_{j}$ at an arbitrary observation time $T_{o}$ is given by
$\mathcal{T}(V)=\begin{cases}&\text{if}\;\Sigma_{r}\geq\Sigma_{T},\;v_{j}=1\\ &\text{if}\;\Sigma_{r}<\Sigma_{T},\;v_{j}=0\\ \end{cases}$ (16)
Determining the state given by equation 15 for every node $v_{j}$ in the network (with node set $V$) by computing equation 15 in parallel at an observation or measurement time $T_{o}$ defines the network dynamical state at $T_{o}$.
Box. 1: Assuming an excitatory network, i.e., a network where for all nodes if a node is activated it generates an excitatory output, we are now in a position to formally restate the key question we ask in this paper: Given some (observed or unobserved) period of evolution dynamics of a network computed by equation 15 for all nodes $v_{j}$ in $G(V,E)$, can the dynamics (state) of the network sustain inherent activity beyond an arbitrary observation time $T_{o}$, or are the dynamics guaranteed to stop, either due to ’epileptic’-like saturation or quiescence? The (global) state of the network is computed by $\mathcal{T}(V)\forall v_{j}\in V$ (equation 16), where $\Sigma_{r}$ is given by equation 15.
@@ 4 Preparing Network States for Deutsch-Jozsa Computations
To leverage the Deutsch–Jozsa algorithm, we need to structure the problem formalized in Box 1. so that it naturally exploits the algorithm’s deterministic superposition properties. This will allow us to determine the global state of the network all at once, rather than needing to (classically) compute and check the state of every $v_{j}$ individually.
To achieve this, we will first construct a specific set that encodes the states of $v_{j}\in V$. When presented as an input to Deutsch–Jozsa, the output can be interpreted as an answer to the question about the network’s dynamics. We will first use Grover’s algorithm to evaluate set membership within this framework.
@@ 4.1 Set Membership Conditions
Each vertex $v_{j}\in V$ is associated with a value $\Sigma_{r}\in\mathbb{R}$ at time $T_{o}$ that is typically normalized to a defined range:
@@
In this context, $\mathcal{T}(V)$ is a function that compares the $n$-bit input $\Sigma_{r}$, the running summation for each vertex $v_{j}$, to a constant threshold $\Sigma_{T}\in\{0,1\}^{n}$ (equation 16). To determine the global state of the network, we evaluate $\mathcal{T}(V)$ for all $v_{j}\in V$ at time $T_{o}$, where the output is either 00 or $1$ for each node, indicating whether $v_{j}$ fires (activates) at the next time step (c.f. Section 3.5.3).
@@
Mathematically, we construct an ordered list $\mathbb{V}_{n}$ whose elements map each vertex $v_{j}\in V$ to its corresponding value of $\Sigma_{r}$:
$\mathbb{V}_{n}:v_{j}\rightarrow\Sigma_{r}\quad\forall v_{j}\in V.$ (17)
Note how this list may contain replicate values of $\Sigma_{r}$.
@@
    $S_{n}\coloneqq\{s_{i}\,|\,s_{i}\in\mathbb{S}_{n}\,\wedge\,s_{i}=\Sigma_{r}\in\mathbb{V}_{n}\}.$ (19)
@@
    $S_{n}^{c}\coloneqq\{s_{i}\,|\,s_{i}\in\mathbb{S}_{n}\,\wedge\,s_{i}\neq\Sigma_{r}\in\mathbb{V}_{n}\}.$ (20)
@@
There exists a surjective mapping $\mathbb{V}_{n}\rightarrow\mathbb{S}_{n}$: Each $v_{j}\in V$ maps to one of the $n$-bit string values $s_{i}\in\mathbb{S}_{n}$. In general, we expect:
$|\mathbb{V}_{n}|=|V|=N>|\mathbb{S}_{n}|\geq|S_{n}|.$
@@
Given this construction, we can use Grover’s algorithm to determine set membership in $S_{n}$. The interpretation of $S_{n}$ in the context of how we framed the neural dynamics question (Box 1.) motivates the construction of the unitary function to be evaluated by the Deutsch–Jozsa algorithm later on and the workflow associated with how we use Deutsch–Jozsa (non-conventionally) so that we can interpret the output of two back to back runs of the algorithm to solve the problem. We progressively develop these ideas over the next few sections.
4.2 Functional Interpretations of $\mathcal{T}(V)$ Outcomes
Recall that our goal is to determine if the network’s dynamics can sustain inherent activity beyond $T_{o}$, or whether it is guaranteed to stop due to either ’epileptic’ saturation or quiescence. This is achieved by evaluating the global state of the network using $\mathcal{T}(V)$ for each node in the network (equation 16). From a computational perspective, given how we constructed $S_{n}$ from $\mathbb{V}_{n}$ (equation 19), the statement in the preceding sentence is equivalent to evaluating $\mathcal{T}(S_{n})\forall s_{i}\in S_{n}$, where as above for $\mathcal{T}(V)$, we will write $\mathcal{T}(s_{i})$ to indicate the evaluation of a specific $s_{i}$ member of $S_{n}$ (see the discussion immediately following equation 16). Formally:
$\mathcal{T}(S_{n})=\begin{cases}&\text{if}\;s_{i}\geq\Sigma_{T},\;s_{i}=1\\ &\text{if}\;s_{i}<\Sigma_{T},\;s_{i}=0\\ \end{cases}$ (21)
Given how we structured the network dynamics question and how the state of the network is computed at the observation $T_{o}$ using equation 16 (Box. 1), we can now also (re)interpret the state of the network by appealing to the evaluation of $\mathcal{T}(S_{n})$:
@@ 4.2 Functional Interpretations of $\mathcal{T}(V)$ Outcomes
    $\mathcal{T}(S_{n})=0\quad\forall s_{i}\in S_{n}$
    In both cases, $\mathcal{T}(S_{n})$ is constant (in the language of Deutsch–Jozsa), mapping $S_{n}$ to binary outputs $\{0,1\}$.
  3. iii.
    In contrast, inherent dynamic activity will occur only when $\mathcal{T}(S_{n})$ is not constant, which we can interpret as the network being able to sustain autonomous activity for at least one additional time step.
The above interpretations allow us to set up a quantum computation whose output resolves the question posed in Box. 1 in the following way:
Box. 2: By running the Deutsch–Jozsa algorithm twice, we can distinguish between constant and non-constant states of $\mathcal{T}(S_{n})$. A non-constant result implies that the network is neither epileptic nor quiescent. It is important to note that $\mathcal{T}(S_{n})$ need not be balanced — only not constant. This agrees with the Deutsch–Jozsa requirement for unitary functions of the form $f(\{0,1\}^{n})\rightarrow\{0,1\}$.
It is critical to note that this is a special use case of the Deutsch-Jozsa algorithm. Deutsch-Jozsa takes as input a function that evaluates a binary sequence of 0’s and 1’s, i.e., a set $\mathbb{S}_{n}=\{0,1\}^{n}$, and uses quantum superposition to simultaneously determine if the function’s output evaluated on the $n$-bit set is constant (all 0’s or all 1’s) or whether the output is balanced (an equal number of 0’s and 1’s). However, in our case here, we need to evaluate $\mathcal{T}(S_{n})$ (equation 21) only on the subset $S_{n}\subset\mathbb{S}_{n}$ using Deutsch-Jozsa (c.f. equation 19). But the algorithm cannot evaluate a particular subset of $\{0,1\}^{n}$. Thus, it cannot explicitly evaluate $\mathcal{T}(S_{n})$ directly. Thus, we first need to extend $S_{n}$ to a construction compatible with the Deutsch-Jozsa, which assumes that the oracle function being evaluated is defined over the entire domain $\mathbb{S}_{n}$. We show how to carry this out below.
@@ 4.3 Determining Membership in $S_{n}$ and $S_{n}^{c}$ Using Grover’s Algorithm
To identify which elements of $\mathbb{S}_{n}$ belong to $S_{n}$ given the dynamics encoded in $\mathbb{V}_{n}$, we need to search for each $n$-bit string $s_{i}\in\mathbb{S}_{n}$ that is in the list $\mathbb{V}_{n}$ which contains the values of $\Sigma_{r}$ resulting from the dynamic evolution of the network up to the observation time $T_{o}$. The resultant set is $S_{n}$, and its complement in $\mathbb{S}_{n}$ forms $S_{n}^{c}$. Recall that we expect some elements of $\mathbb{S}_{n}$ to appear in $\mathbb{V}_{n}$ only once, some multiple times, and others not at all. Grover’s algorithm offers a potential quadratic speedup over classical approaches to construct $S_{n}$ by enabling an efficient search within $\mathbb{S}_{n}$. In particular, for our purposes here, we are effectively using Grover to search for the values in $\mathbb{S}_{n}$ that belong to $S_{n}$ in order to ‘mark’ them with a phase-flip in a quantum circuit, and not necessarily to amplify them. This will be necessary because the members of $S_{n}$ will need to be functionally evaluated by equation 21 using a bit-subtraction approach outlined below. The elements of the complement $S_{n}^{c}$ will not be evaluated by equation 21 and will instead be assigned a constant value of 00 or $1$ for each of the two Deutsch-Josza runs. The practical reasons for leveraging Grover for this task become evident when attempting to implement this model in a series of quantum circuits. We return to this below in Section 7.2.
@@ 4.3.1 Construction of the oracle function
The list $\mathbb{V}_{n}$ contains the running summation values $\Sigma_{r}$ for each vertex $v_{j}$ at $T_{o}$ (see equation 17). Assume each element in $\mathbb{V}_{n}$ is labeled as $\{\nu_{1},\nu_{2},\dots,\nu_{M}\}$. Each of these elements of $\mathbb{V}_{n}$ corresponds to an $n$-bit string in $\mathbb{S}_{n}$. For each $s_{i}\in\mathbb{S}_{n}$, we construct an oracle function $U_{f}$ to identify if $s_{i}\in\mathbb{V}_{n}$. Specifically, $U_{f}$ checks if there exists a $\nu_{i}\in\mathbb{V}_{n}$ such that $\nu_{i}=s_{i}$ for some $s_{i}\in\mathbb{S}_{n}$:
$\exists\nu_{i}\in\mathbb{V}_{n}\text{ such that }\nu_{i}=s_{i}.$ (22)
@@
where |$\ket{\nu_{i}}$ encodes a state corresponding to $\nu_{i}\in\mathbb{V}_{n}$, and $f(\nu_{i})$ evaluates to 1 if $\nu_{i}=s_{i}$, and 0 otherwise.
To construct $U_{f}$, we will use a standard gate-based construction of Grover that implements the logical condition ‘if the current state $\nu_{i}=s_{i}$, apply a phase flip; otherwise, do nothing’(Section 7.2).